# Can Large Audio-Language Models Truly Hear? Tackling Hallucinations with Multi-Task Assessment and Stepwise Audio Reasoning

### v1

@@ I Introduction
To explore model hallucinations, we develop tasks addressing object existence, temporal order, and object attributes, as illustrated in Fig. 1. For more examples and details, please refer to the demo site111 https://sites.google.com/view/lalm-hallucination-icassp2025 . Our dataset consists of 10,800, 3,116, and 1,614 instances for object existence, temporal order, and object attributes, respectively. The object existence task evaluates the models’ ability to detect specific sound events. The temporal order task assesses the models’ capability to determine the sequence of sound events. The object attribute task examines the models’ skill in identifying sound sources. We formulated these tasks as discriminative questions[37], such as asking about the presence of a dog barking or the relative timing of tiger and human sounds. These binary classification questions were evaluated using standard metrics such as accuracy, precision, recall, and F1 score. To extend this approach, we introduced paired question sets inspired by previous research[38]. These pairs consist of questions about two closely related audio segments: a “before” segment and an “after” segment. The “after” segment is identical to the “before” segment, except for the deliberate removal of one sound event. This design assesses the models’ ability to detect changes in sound events across sequential audio clips. For instance, in object existence section of Fig 1, a “before” clip might contain dog barking, cat meowing, and someone coughing, while the corresponding “after” clip would only include cat meowing and someone coughing. By comparing the models’ responses to questions about both segments, we can evaluate its sensitivity to changes in audio content over time.
@@ II Method
Qwen-Audio-Chat 49.8 52.1 15.4 23.7 10.2 75.7 84.0 98.5 49.1 49.7 38.3 43.3 18.3 43.2 47.5 55.9 16.4 25.4 11.0 76.0 49.7 50.4 35.7 41.8 18.0 47.2
Qwen2-Audio-Instruct 50.1 62.5 0.6 1.3 0.5 99.1 99.4 99.9 50.4 59.4 2.4 4.7 1.8 96.5 50.4 58.3 2.7 5.2 1.5 96.5 52.2 54.2 28.4 37.3 17.5 58.6
SALMONN-7B 50.0 50.7 2.4 4.5 1.9 95.8 97.7 100.0 50.2 52.7 3.7 7.0 3.0 93.7 50.0 50.0 0.3 0.5 0.3 99.5 49.9 49.9 14.9 22.9 10.4 74.7
SALMONN-13B 51.1 53.0 19.3 28.3 11.7 71.3 81.8 100.0 50.7 51.9 20.4 29.2 12.1 69.0 50.0 51.9 0.9 1.8 0.7 98.5 54.0 52.6 81.7 64.0 19.3 7.1
LTU-AS 49.3 49.3 43.5 46.2 15.8 39.5 55.8 99.9 49.3 49.3 54.2 51.6 15.0 29.3 49.5 49.4 39.4 43.8 15.9 43.7 50.6 50.9 42.5 46.3 18.4 40.5
Gazelle 48.1 50.4 33.0 39.9 10.6 56.2 64.7 96.8 49.0 49.6 30.5 37.7 11.2 57.6 51.5 52.1 40.2 45.3 17.4 45.8 48.5 50.3 83.6 62.8 12.3 5.5
GAMA 49.8 20.0 0.1 0.1 0.1 99.7 99.7 99.8 50.0 51.9 0.9 1.8 0.7 98.5 49.4 40.0 1.4 2.8 0.9 97.3 50.5 51.0 25.1 33.7 12.7 63.3
Gemini-1.5-flash 43.7 49.6 24.5 32.8 12.1 63.3 66.0 87.5 40.1 49.9 67.9 57.6 11.1 19.0 17.9 50.3 73.7 59.8 8.0 18.3 42.6 50.4 92.0 65.1 10.2 0.6
Gemini-1.5-pro 35.5 52.5 65.7 58.4 18.9 19.7 24.6 67.9 39.7 50.6 85.3 63.5 9.9 6.0 25.7 50.5 93.7 65.6 3.7 1.1 43.2 50.4 90.8 64.8 10.1 1.2
Cascade 41.1 50.6 92.2 63.8 5.2 2.4 6.3 81.3 37.1 50.4 89.7 64.5 4.9 3.2 31.4 51.0 87.6 64.5 6.0 5.4 - - - - - -

### v2

@@ I Introduction
To explore model hallucinations, we develop tasks addressing object existence, temporal order, and object attributes, as illustrated in Fig. 1. For more examples, code, and benchmarks, visit our project page111 github.com/kuan2jiu99/audio-hallucination. Our dataset consists of 10,800, 3,116, and 1,614 instances for object existence, temporal order, and object attributes, respectively. The object existence task evaluates the models’ ability to detect specific sound events. The temporal order task assesses the models’ capability to determine the sequence of sound events. The object attribute task examines the models’ skill in identifying sound sources. We formulated these tasks as discriminative questions[37], such as asking about the presence of a dog barking or the relative timing of tiger and human sounds. These binary classification questions were evaluated using standard metrics such as accuracy, precision, recall, and F1 score. To extend this approach, we introduced paired question sets inspired by previous research[38]. These pairs consist of questions about two closely related audio segments: a “before” segment and an “after” segment. The “after” segment is identical to the “before” segment, except for the deliberate removal of one sound event. This design assesses the models’ ability to detect changes in sound events across sequential audio clips. For instance, in object existence section of Fig 1, a “before” clip might contain dog barking, cat meowing, and someone coughing, while the corresponding “after” clip would only include cat meowing and someone coughing. By comparing the models’ responses to questions about both segments, we can evaluate its sensitivity to changes in audio content over time.
@@ II Method
Qwen-Audio-Chat 49.5 51.2 15.1 23.3 9.3 75.4 84.0 98.5 48.8 49.8 38.1 43.2 18.1 43.7 47.4 54.7 16.1 24.9 11.7 75.8 48.2 48.5 34.3 40.2 16.9 48.2
Qwen2-Audio-Instruct 50.0 56.2 0.6 1.1 0.4 99.2 99.4 99.9 50.6 65.2 2.8 5.3 2.3 96.1 50.3 55.6 2.6 4.9 1.6 96.7 52.7 55.2 28.9 37.9 18.6 58.5
SALMONN-7B 50.0 49.3 2.3 4.4 1.7 96.2 97.7 100.0 50.6 58.2 4.1 7.7 2.9 94.4 49.9 37.5 0.2 0.4 0.3 99.4 52.5 58.7 17.5 27.0 11.3 76.5
SALMONN-13B 53.4 59.4 21.6 31.6 12.6 74.6 81.8 100.0 53.6 59.2 23.2 33.4 13.3 72.1 50.4 70.4 1.2 2.4 0.8 98.9 58.1 55.2 85.8 67.2 23.9 8.5
LTU-AS 50.5 50.6 44.7 47.5 16.7 37.2 55.8 99.9 50.4 50.3 55.3 52.7 15.7 27.1 49.4 49.2 39.3 43.7 15.9 41.9 51.3 51.8 43.2 47.1 19.0 40.1
Gazelle 47.9 49.7 32.8 39.5 9.9 58.4 64.7 96.8 49.0 49.5 30.4 37.7 11.2 56.9 50.4 50.5 39.0 44.0 17.3 46.8 48.7 50.6 83.6 63.1 13.2 6.3
GAMA 49.9 40.0 0.1 0.3 0.1 99.7 99.7 99.8 50.1 59.3 1.0 2.0 0.9 98.3 49.3 40.0 1.4 2.7 0.8 97.1 50.3 50.7 24.9 33.4 12.0 66.0
Gemini-1.5-flash 43.5 49.2 24.2 32.5 12.1 62.4 66.0 87.5 39.5 48.8 67.5 56.7 10.4 18.0 18.0 49.9 74.8 59.9 8.3 18.3 44.0 51.0 95.6 66.5 11.6 0.4
Gemini-1.5-pro 35.7 52.7 66.0 58.6 19.9 18.1 24.6 67.9 40.1 50.7 86.0 63.8 9.8 5.5 26.0 50.5 94.7 65.9 3.7 1.5 44.9 51.5 94.3 66.6 10.7 0.7
Cascade 40.6 50.3 91.7 65.0 5.5 2.5 6.3 81.3 37.7 51.3 89.9 65.4 5.0 3.3 31.5 51.7 87.2 64.9 5.8 5.4 - - - - - -
@@ V Conclusions
VI Acknowledgement
We would like to thank Wei-Ping Huang and Hsuan Su for providing valuable feedback on the draft of this paper. Additionally, we thank the National Center for High-performance Computing (NCHC) of National Applied Research Laboratories (NARLabs) in Taiwan for providing computational and storage resources.